# PointCFormer: a Relation-based Progressive Feature Extraction Network for Point Cloud Completion

### v1

@@ Abstract
Point cloud completion aims to reconstruct the complete 3D shape from incomplete point clouds, and it is crucial for tasks such as 3D object detection and segmentation. Despite the continuous advances in point cloud analysis techniques, feature extraction methods are still confronted with apparent limitations. The sparse sampling of point clouds, used as inputs in most methods, often results in a certain loss of global structure information. Meanwhile, traditional local feature extraction methods usually struggle to capture the intricate geometric details. To overcome these drawbacks, we introduce PointCFormer, a transformer framework optimized for robust global retention and precise local detail capture in point cloud completion. This framework embraces several key advantages. First, we propose a relation-based local feature extraction method to perceive local delicate geometry characteristics. This approach establishes a fine-grained relationship metric between the target point and its k-nearest neighbors, quantifying each neighboring point’s contribution to the target point’s local features. Secondly, we introduce a progressive feature extractor that integrates our local feature perception method with self-attention. Starting with a denser sampling of points as input, it iteratively queries long-distance global dependencies and local neighborhood relationships. This extractor maintains enhanced global structure and refined local details, without generating substantial computational overhead. Additionally, we develop a correction module after generating point proxies in the latent space to reintroduce denser information from the input points, enhancing the representation capability of the point proxies. PointCFormer demonstrates state-of-the-art performance on several widely used benchmarks.
@@ Introduction
Among these existing methods, the feature extraction of points is an essential step, which primarily relies on the combined use of down-sampling and k-nearest neighbor (k-NN) search. However, the risk of losing certain global structural information is inevitable if an overly sparse down-sampled point cloud is employed as the model input. In addition, traditional feature extraction algorithms based on simple k-NN search can result in the network only being able to extract relatively coarse local features. Unfortunately, these fundamental limitations have seldom been discussed in previous works.
In this paper, we delve into the utilization of an advanced feature extractor for point cloud completion, aiming to mitigate the loss of global information during point cloud sampling and the coarse extraction issues in the local feature extraction process. We propose a framework, termed PointCFormer, as shown in Fig. 2. Specifically, PointCFormer adopts the transformer-based encoder-decoder architecture. To handle the coarse extraction problem, we propose fine-grained computation for the contribution of the neighboring point to the target point according to the geometry characteristics of the local region (including three-dimensional spatial coordinates and high-dimensional feature vectors). To better preserve the global shape information during the sampling process, we introduce a progressive method with slow down-sampling, which gradually refines the features through alternately global and local queries. Compared to previous work, Fig. 1 demonstrates the superior performance of our method in the point cloud completion task in terms of both global shape and local details. In summary, our main contributions are as follows:
@@
For 3D point cloud completion, traditional approaches (Dai, Ruizhongtai Qi, and Nießner 2017; Han et al. 2017; Stutz and Geiger 2018) employ voxelization or distance fields to describe 3D objects and process them using 3D convolutional neural networks (Wu et al. 2015). However, the high demand for memory and computational resources limits their application scope (Wang et al. 2022). To mitigate this issue, researchers have shifted towards using unstructured point clouds to represent 3D objects, exploring various innovative approaches. PointNet and its variants (Qi et al. 2017a, b) extract features directly from unstructured point clouds, provide a new perspective on the task of 3D point cloud processing. PCN (Yuan et al. 2018) leverages an encoder-decoder architecture and simulates the deformation process of the 2D plane through the FoldingNet technique (Yang et al. 2018). This maps 2D points onto 3D surfaces to achieve point cloud completion while preserving geometric structure and topological properties. SnowflakeNet (Xiang et al. 2021) models the point cloud generation process as a snowflake-like growth pattern based on certain base points in 3D space. This pattern gradually expands from the parent point to form a complete 3D shape, simulating complex geometric structures. LAKeNet (Tang et al. 2022) introduces a novel 3D shape prediction method, which effectively captures the key features of 3D shapes by considering the structured and topological information of the predicted 3D shapes and following a key point-skeleton-shape prediction process. PointAttN (Wang et al. 2024) transforms the traditional k-NN method of obtaining local features into an attention-based method. This refines the global dependency on the relationships between point clouds. SeedFormer (Zhou et al. 2022) introduces patch seeds as a new shape representation for point clouds and designs an upsampling transformer to make point cloud completion more efficient and accurate. Anchorformer (Chen et al. 2023) enhances the accuracy of point cloud completion by capturing region information through pattern recognition nodes and combining the information of sampling points and anchor points. Recently, 3DMambaComplete  (Li, Yang, and Fei 2024) applies the Mamba (Gu and Dao 2023; Dao and Gu 2024) structure to point cloud completion, showcasing its advantages in processing 3D point cloud data, which brings new insights to the community.
@@
Figure 3: Comparison of local feature extraction: traditional kNN-based method (Left); DGCNN with relative relation fusion (Center); Ours (Right).
@@
The k-NN search may identify neighbors with inaccurate relevance to the target points due to sparse sampling, as illustrated in Fig. 4(a), where only the two closest points (orange) show significant correlation with the target points (red), and more distant points (green) may provide little or even misleading information. To solve this problem, we propose a scheme ( Fig. 3(c)) that introduces an adaptative contribution weight for each nearest neighbor based on its correlation with the target point. The redder the line segments in Figure  3(c) the stronger the correlation between the two points connected and the bluer the line segments the weaker the correlation between the two points connected. Specifically, we quantify the interrelationships between a target point and its neighboring points in terms of geometry in 3D space and feature vectors in high-dimensional space simultaneously. The quantified values are then utilized in a learning-based approach to assess each neighbor’s contribution to the target point.
@@
$R_{2}\left(Q,V_{i}\right)=\left|\frac{1}{M}\sum_{j=1}^{M}\operatorname{vec}\left(Q,V_{j}\right)-\operatorname{vec}\left(Q,V_{i}\right)\right|$ (2)
@@
Fig. 5 illustrates the technical implementation of local geometric relationship perception. Specifically, the left part utilizes the process of EdgeConv for local feature extraction. The right part is our proposed relationship perception. We concatenate the relationship metrics $R1$ and $R2$ and use an MLP to output contribution weights in a learning manner. Then, the generated weights are element-wise multiplied by the extracted features. Finally, robust and geometry-related features are generated through a max function. We provide a more detailed analysis of local geometric perception with respect to the relationship metrics $R1$ and $R2$ in the Supplementary Material.
@@
In this subsection, PointCFormer is compared with several advanced methods, including FoldingNet (Yang et al. 2018), PCN (Yuan et al. 2018), GRNet (Xie et al. 2020), PoinTr (Yu et al. 2021), SnowFlakeNet (Xiang et al. 2021), SeedFormer (Zhou et al. 2022), AnchorFormer (Chen et al. 2023), PointAttN (Wang et al. 2024), and AdaPointr (Yu et al. 2023).
@@
The PCN dataset comprises 28,974 shapes for training and 1,200 shapes for testing, distributed across eight categories. It is currently the most frequently utilized benchmark dataset for evaluating the performance of point cloud completion methods. As shown in Table 1, our method leads in almost all metrics. Specifically, our method improves the average CD-$\ell_{1}$ by 0.43 compared to the recently proposed PointAttN and shows remarkable improvements across different categories. Compared to the 3DMambaCom, our average CD-$\ell_{1}$ improves by 0.5, and we see a 0.31 increase in the F-Score@1%. As illustrated in Table 2, our methodology, while drawing elements from AdaPoinTr, is not simply an incremental model. Instead, it shows marked improvement in overall performance. This is largely due to our innovative relation-based local feature perception and the progressive optimization in feature extraction.
@@
Table 2: Further comparison of our method with AdaPoinTr and its expanded size version.
Method Params FLOPs Avg CD-$\ell_{1}$ F-Score@1%
AdaPoinTr 32.4M 241G 6.54 0.841
AdaPoinTr-Large 39.6M 271G 6.49 0.847
PointCFormer 33.2M 274G 6.41 0.855
@@
Next, we evaluate PointCFormer on the ShapeNet-55 dataset, which encompasses a wider range of categories. Table 3 presents the performance of various methods on incomplete point cloud data with three different missing ratios (CD-$\ell_{2}$-S, CD-$\ell_{2}$-M, and CD-$\ell_{2}$-H) in terms of L2 Chamfer distance (CD-$\ell_{2}$) and F-Score@1%. PointCFormer evidently outperforms other existing methods in CD-$\ell_{2}$-S, CD-$\ell_{2}$-M, CD-$\ell_{2}$-H, and average CD-$\ell_{2}$. Despite scoring slightly lower than the top-performing method (AdaPoinTr) in terms of F-Score@1%, we achieve state-of-the-art performance with a CD-$\ell_{2}$ of 0.73. These results underscore the effective completion capabilities of PointCFormer, even when applied to such a diverse dataset.
@@
More visual comparisons on several benchmarks are shown in the Supplementary Material.
@@
To demonstrate the effectiveness of each module we designed within PointCFormer, we conduct a comprehensive ablation study on the PCN dataset. The results are summarized in Table 6. Here, “LGRP” refers to a progressive feature extractor that contains only two ST-LFE modules; “GSP” refers to a progressive feature extractor that uses only two self-attention layers (Applying LGRP and GSP is the whole PFE.), and “CM” stands for the correction module. The baseline “I” is a combination of the partial framework from AdaPoinTr with a conventional encoder. We observe a significant performance boost when we replace the traditional local feature extraction process with “LGRP”. The model’s performance was further enhanced upon the incorporation of the global shape perception process. Lastly, the best performance of the model was obtained after adding the correction module. It is evident from the results that the local geometric relationship perception contributes the most to the model’s improvement, showcasing its great potential for other point cloud tasks.
@@ Conclusions
In this paper, we explore preserving extensive global structural features while extracting detailed local features in point cloud completion tasks. We introduce PointCFormer, a novel point cloud completion model that builds upon the Transformer framework. A key component is the local relation perception module, which measures the contribution of each point within a local region based on the relations in 3D space and high-dimensional feature space. Another key component is the progressive feature extractor combined with the attention mechanism to strengthen the capture of global structures. Additionally, the latent space correction module ensures that the generated point proxies align more accurately with the distribution of the original input point cloud data. Crucially, we incorporate a relation-based local perception mechanism into both the feature extractor and the correction module, significantly enhancing their performance. Comprehensive comparative and ablation studies across various challenging benchmarks demonstrate that PointCFormer achieves superior completion performance. In fact, our relation-based feature extraction method comprehensively considers the global and local information of point clouds. This method can be naturally extended to other point cloud-related tasks, which will be explored and evaluated in the future.

### v2

@@ Abstract
Point cloud completion aims to reconstruct the complete 3D shape from incomplete point clouds, and it is crucial for tasks such as 3D object detection and segmentation. Despite the continuous advances in point cloud analysis techniques, feature extraction methods are still confronted with apparent limitations. The sparse sampling of point clouds, used as inputs in most methods, often results in a certain loss of global structure information. Meanwhile, traditional local feature extraction methods usually struggle to capture the intricate geometric details. To overcome these drawbacks, we introduce PointCFormer, a transformer framework optimized for robust global retention and precise local detail capture in point cloud completion. This framework embraces several key advantages. First, we propose a relation-based local feature extraction method to perceive local delicate geometry characteristics. This approach establishes a fine-grained relationship metric between the target point and its k-nearest neighbors, quantifying each neighboring point’s contribution to the target point’s local features. Secondly, we introduce a progressive feature extractor that integrates our local feature perception method with self-attention. Starting with a denser sampling of points as input, it iteratively queries long-distance global dependencies and local neighborhood relationships. This extractor maintains enhanced global structure and refined local details, without generating substantial computational overhead. Additionally, we develop a correction module after generating point proxies in the latent space to reintroduce denser information from the input points, enhancing the representation capability of the point proxies. PointCFormer demonstrates state-of-the-art performance on several widely used benchmarks. Our code is available at https://github.com/Zyyyyy0926/PointCFormer_Plus_Pytorch.
@@ Introduction
Among these existing methods, the feature extraction of points is an essential step, which primarily relies on the combined use of down-sampling and k-nearest neighbor (k-NN) search. However, the risk of losing certain global structural information is inevitable if an overly sparse down-sampled point cloud is employed as the model input. In addition, traditional feature extraction algorithms based on simple k-NN search can result in the network only extracting relatively coarse local features. Unfortunately, these fundamental limitations have seldom been discussed in previous works.
In this paper, we delve into the utilization of an advanced feature extractor for point cloud completion, aiming to mitigate the loss of global information during point cloud sampling and the coarse extraction issues in local feature extraction. We propose a framework, termed PointCFormer, as shown in Fig. 2. Specifically, PointCFormer adopts the transformer-based encoder-decoder architecture. To handle the coarse extraction problem, we propose fine-grained computation for the contribution of the neighboring point to the target point according to the geometry characteristics of the local region (including 3D spatial coordinates and high-dimensional feature vectors). To better preserve the global shape information during the sampling process, we introduce a progressive method with slow down-sampling, which gradually refines the features through alternately global and local queries. Compared to previous work, Fig. 1 shows the superior performance of our method in the point cloud completion task in terms of global shape and local details. In summary, our main contributions are as follows:
@@
For 3D point cloud completion, traditional approaches (Dai, Ruizhongtai Qi, and Nießner 2017; Han et al. 2017; Stutz and Geiger 2018) employ voxelization or distance fields to describe 3D objects and process them using 3D convolutional neural networks (Wu et al. 2015). However, the high demand for memory and computational resources limits their application scope (Wang et al. 2022). To mitigate this issue, researchers have shifted towards using unstructured point clouds to represent 3D objects, exploring various innovative approaches. PointNet and its variants (Qi et al. 2017a, b) extract features directly from unstructured point clouds, provide a new perspective on the task of 3D point cloud processing. PCN (Yuan et al. 2018) leverages an encoder-decoder architecture and simulates the deformation process of the 2D plane through the FoldingNet technique (Yang et al. 2018). This maps 2D points onto 3D surfaces to achieve point cloud completion while preserving geometric structure and topological properties. SnowflakeNet (Xiang et al. 2021) models the point cloud generation process as a snowflake-like growth pattern based on certain base points in 3D space. This pattern gradually expands from the parent point to form a complete 3D shape, simulating complex geometric structures. LAKeNet (Tang et al. 2022) introduces a novel 3D shape prediction method, which effectively captures the key features of 3D shapes by considering the structured and topological information of the predicted 3D shapes and following a key point-skeleton-shape prediction process. PointAttN (Wang et al. 2024) transforms the traditional k-NN method of obtaining local features into an attention-based method. This refines the global dependency on the relationships between point clouds. SeedFormer (Zhou et al. 2022) introduces patch seeds as a new shape representation for point clouds and designs an upsampling transformer to make point cloud completion more efficient and accurate. Anchorformer (Chen et al. 2023) enhances the accuracy of point cloud completion by capturing region information through pattern recognition nodes and combining the information of sampling points and anchor points. Recently, (Li, Yang, and Fei 2024) applies the Mamba (Gu and Dao 2023) structure to point cloud completion, showcasing its advantages in processing 3D point cloud data, which brings new insights to the community.
@@
Figure 3: Comparison of local feature extraction: traditional kNN-based method (Left); DGCNN with relative relation fusion (Center); Ours (Right). In (c), the redder the line segment, the stronger the correlation between the two connected points; the bluer the line segment, the weaker the correlation.
@@
The k-NN search may identify neighbors with inaccurate relevance to the target points due to sparse sampling, as illustrated in Fig. 4(a), where only the two closest points (orange) show significant correlation with the target points (red), and more distant points (green) may provide little or even misleading information. To solve this problem, we propose a scheme (Fig. 3(c)) that introduces an adaptative contribution weight for each nearest neighbor based on its correlation with the target point. Specifically, we quantify the interrelationships between a target point and its neighboring points in terms of geometry in 3D space and feature vectors in high-dimensional space simultaneously. The quantified values are then utilized in a learning-based approach to assess each neighbor’s contribution to the target point.
@@
$R_{2}\left(Q,V_{i}\right)=\left|\frac{1}{M}\sum_{j=1}^{M}\operatorname{vec}\left(Q,V_{j}\right)-\operatorname{vec}\left(Q,V_{i}\right)\right|,$ (2)
@@
Fig. 5 illustrates the technical implementation of local geometric relationship perception. Specifically, the left part utilizes the process of EdgeConv for local feature extraction. The right part is our proposed relationship perception. We concatenate the relationship metrics $R1$ and $R2$ and use an MLP to output contribution weights in a learning manner. Then, the generated weights are element-wise multiplied by the extracted features. Finally, robust and geometry-related features are generated through a max function. We provide a more detailed analysis of local geometric perception with respect to the relationship metrics $R1$ and $R2$ in the Supplementary Material(Zhong et al. 2024).
@@
We compare our method with several advanced methods, including FoldingNet (Yang et al. 2018), PCN (Yuan et al. 2018), GRNet (Xie et al. 2020), PoinTr (Yu et al. 2021), SnowFlakeNet (Xiang et al. 2021), SeedFormer (Zhou et al. 2022), AnchorFormer (Chen et al. 2023), PointAttN (Wang et al. 2024), and AdaPointr (Yu et al. 2023).
@@
The PCN dataset comprises 28,974 shapes for training and 1,200 shapes for testing, distributed across eight categories. It is currently the most frequently utilized benchmark dataset for evaluating the performance of point cloud completion methods. As shown in Table 1, our method leads in almost all metrics. Specifically, our method improves the average CD-$\ell_{1}$ by 0.43 compared to the recently proposed PointAttN and shows remarkable improvements across different categories. Compared to the 3DMambaCom, our average CD-$\ell_{1}$ improves by 0.5, and we see a 0.31 increase in the F-Score@1%.
@@
Next, we evaluate PointCFormer on the ShapeNet-55 dataset, which encompasses more categories. Table 2 presents the performance of various methods on incomplete point cloud data with three different missing ratios (CD-$\ell_{2}$-S, CD-$\ell_{2}$-M, and CD-$\ell_{2}$-H) in terms of L2 Chamfer distance (CD-$\ell_{2}$) and F-Score@1%. PointCFormer evidently outperforms other existing methods in CD-$\ell_{2}$-S, CD-$\ell_{2}$-M, CD-$\ell_{2}$-H, and average CD-$\ell_{2}$. Despite scoring slightly lower than the top-performing method (AdaPoinTr) in terms of F-Score@1%, we achieve state-of-the-art performance with a CD-$\ell_{2}$ of 0.73. These results underscore the effective completion capabilities of PointCFormer, even when applied to such a diverse dataset.
@@
We conducted a comprehensive ablation study on the PCN dataset to demonstrate the effectiveness of each module we designed within PointCFormer. The results are summarized in Table 5. Here, “LGRP” refers to a progressive feature extractor that contains only two ST-LFE modules, “GSP” refers to a progressive feature extractor that uses only two self-attention layers (Applying LGRP and GSP is the whole PFE.), and “CM” stands for the correction module. The baseline “I” is a combination of the partial framework from AdaPoinTr with a conventional encoder. We observe a significant performance boost when we replace the traditional local feature extraction process with “LGRP”. The model’s performance was further enhanced by incorporating the global shape perception process (GSP). Lastly, the model’s best performance was obtained after adding the correction module. It is evident from the results that the local geometric relationship perception contributes the most to the model’s improvement, showcasing its great potential for other point cloud tasks.
@@ Conclusions
In this paper, we explore preserving extensive global structural features while extracting detailed local features in point cloud completion tasks. We introduce PointCFormer, a novel point cloud completion model that builds upon the Transformer framework. A key component is the local relation perception module, which measures the contribution of each point within a local region based on the relations in 3D space and high-dimensional feature space. Another key component is the progressive feature extractor combined with the attention mechanism to strengthen the capture of global structures. Additionally, the latent space correction module ensures that the generated point proxies align more accurately with the distribution of the original input point cloud data. Crucially, we incorporate a relation-based local perception mechanism into both the feature extractor and the correction module, significantly enhancing their performance. Comprehensive comparative and ablation studies across various challenging benchmarks demonstrate that PointCFormer achieves superior completion performance. In fact, our relation-based feature extraction method comprehensively considers the global and local information of point clouds. We would like to extend our method to other point cloud-related tasks in the future.
Acknowledgments
This work is partially funded by the National Natural Science Foundation of China (12494550 and 12494553), the Beijing Natural Science Foundation (Z240002), and the Guangdong Science and Technology Program (2023B1515120026).